GDP= 100,000, taxes = 22,000, government purchases = 25,000 and national savings =15,000. For this economy , consumption amounts to

a) 38,000

b) 60,000

c) 68,000

d) 53,000
Answer #1

Answer

the economy is closed as there is no import or export so the equation of GDP is:

GDP=C+I+G

where

S=saving=I=investment

G=25000

I=15000

GDP=100000

Consumption =GDP-G-I

=100000-25000-15000

=$60000

the consumption is $60000

option b